# Quantum critical theories in a periodic potential: strange metallic thermoelectric and magneto-transport

### v1

@@ 3 Transport expectations from hydrodynamics
The periodic potential environment in which we wish to study the quantum critical response, is built in the UV boundary conditions of our ansatz 10. At finite density, i.e., finite average $\bar{\mu}\neq 0$, a simple (relativistic) hydrodynamic analysis extended with a momentum relaxation rate $\Gamma$ parametrically smaller than any other rate ($\bar{\mu}\gg\delta\mu$ in the case of an chemical potential lattice) results in thermo-electric transport that is dominated by a Drude pole at $\omega=-i\Gamma\sim-i(\delta\mu)^{2}$, such that the AC transport coefficients take the general form davison_dissecting_2015; hartnoll_theory_2007
@@
where $n,s,\varepsilon$ and $p$ denote the charge, entropy, energy and pressure densities, respectively. We will be studying the system in the opposite limit $\bar{\mu}\ll\delta\mu$, but the generic wisdom that translational symmetry breaking leads to finite DC conductivities should hold. In the extremal limit of interest to us, $\bar{\mu}=n=0$, a simple exercise in hydrodynamics indicates that electric and thermal transport will decouple. We might therefore expect the electrical conductivity to be governed by the incoherent conductivity $\sigma_{Q}$. Similarly, the thermal conductivity should remain coherent (i.e., dominated by a single pole). In the absence of magnetic fields, we furthermore expect $\alpha=0$, confirming that the mechanisms of electrical and thermal transport are essentially independent. We thus also expect no difference between the thermal conductivity at zero electric current $\kappa_{ij}=\bar{\kappa}_{ij}-T\bar{\alpha}_{ij}(\sigma^{-1})_{kl}\alpha_{lj}$} and $\bar{\kappa}_{ij}$.
At finite magnetic fields, the transport coefficients become tensorial in nature, and one naturally expects $\sigma_{xy}\sim n/B$, as follows from a combination of parity and charge-conjugation symmetry. In a charge neutral system with $\bar{n}=0$, however, no Hall effect will occur, and the magnetic field will only affect longitudinal transport. Nevertheless, a finite magnetic field does allow for a Hall thermoelectric conductivity $\alpha_{xy}\sim s/B$ and we will be able to observe a non-zero contribution to the Nernst coefficient $e_{N}=-(\sigma^{-1}\alpha)_{xy}$, where from charged hydrodynamics one expects hartnoll_theory_2007
@@
with the cyclotron frequency in natural units equal to $\omega_{c}=\frac{Bn}{\epsilon+p}$. For a charge neutral system this simplifies to a direct measurement of the momentum relaxation rate as $e_{N}=\frac{B}{T\Gamma}$.
Figure 1: DC quantum critical transport in a 1D charge-neutral lattice, $\mu(x)=\delta\mu\cos(Gx)$}. Note that the thermoelectric conductivity $\alpha=0$} (numerically confirmed, not shown) and thus $\bar{\kappa}=\kappa$. (a) Electrical conductivity $\sigma$ as a function of $T/\delta\mu$ for various $\delta\mu/G$ (left colorbar). At low temperatures, the lattice causes the conductivity to increase. Inset: electrical conductivity as a function of $T/G$. (b) Electrical conductivity $\sigma$ as a function of lattice amplitude for various temperatures $T/G$ (right colorbar). In the low-temperature limit for small $\delta\mu/G$, the scaling is that predicted by Chesler, Lucas and Sachdev chesler_conformal_2014, (Eq. 28 black dashed line). Inset: difference from the CFT value $\sigma_{\infty}=1$ on a logarithmic scale, which grows as $(\delta\mu)^{2}$. (c) Thermal conductivity $\bar{\kappa}$ as a function of $T/\delta\mu$. There is a cross-over from a universal high-$T$ regime to a strong thermal conductor-regime at low $T$. Inset: thermal conductivity $\bar{\kappa}$ as a function of $T/G$. (d) Thermal conductivity $\bar{\kappa}$ as a function of lattice amplitude. The thermal conductivity decreases most rapidly with increased lattice amplitude around $T/G\approx 0.1$}, which corresponds with the minima seen in panel (c). Upper inset: for weak lattices, the thermal conductivity scales as $1/(\delta\mu)^{2}$. Lower inset: for strong lattices, the thermal conductivity approaches the bound $\kappa/T\geq 4\pi^{2}/3$ of grozdanov_incoherent_2016.
There is one more expectation we can formulate. Our system is on average isotropic in the thermodynamic limit. In quantum critical theories (with a holographic dual), this implies that the DC electrical conductivity is bounded from below as $\sigma_{DC}\geq\sigma_{\infty}$ grozdanov_absence_2015. The electrical conductivity at charge neutrality in the absence of a lattice is given by that of the CFT dual to the Schwarzchild AdS4 solution: $\sigma_{DC}=\sigma_{\infty}=1$. This saturates the bound. Thus, the only thing the conductivity can do is to increase with increased disorder/increased lattice amplitude around $\bar{\mu}=0$. We will indeed see this below.
@@ 4.1 DC transport in 1D charge-neutral lattices
To build on the results of Chesler, Lucas and Sachdev chesler_conformal_2014, we first consider a 1D lattice at charge neutrality (i.e., that is charge-conjugation symmetric): we set the background chemical potential to be
@@
This scaling of the electrical conductivity $\sigma_{DC}$ as a function of lattice strength was elucidated by Chesler, Lucas and Sachdev chesler_conformal_2014: deforming the UV with a 1D lattice with $\bar{\mu}=0$ still results in an AdS4 IR geometry at $T=0$, except that in the natural UV coordinates the IR metric is anisotropic with regards to the $x/y$-directions. We highlight this effect in Figure 2 (left column), that shows the RG flow (captured by the radial coordinate $z$), but halted due to the finite value of $T$, of the density vanishing in the IR and two metric components exhibiting the anisotropy. Absorbing the emergent anisotropy in the coordinates to bring the IR metric into standard AdS4 form also rescales the length scales of the dual theory, resulting in
@@
This perturbative result, valid for $T=0$ and at $\delta\mu\ll G$, is shown shown as a black dashed line in Figure 1(b) and which can be seen to asymptotically hold for $T\ll G$; in fact, the result extends to $\delta\mu/G=O(1)$ for low enough temperatures. At larger temperatures the deviation from $\sigma_{\infty}$ is not as strong, but still scales as $(\delta\mu)^{2}$ (inset, panel (b)). Two comments about this strong coordinate anisotropy hiding translational symmetry mechanism are in order. Firstly, the explanation is particular to 1D, and will not apply to the 2D lattice as we shall see in Section 5. Secondly, as explained by Chesler, Lucas and Sachdev, this emergent translational symmetry is an artifact of the holographic large $N$ approximation chesler_conformal_2014. There it was shown that by comparing to a system of $N_{f}$ free Dirac fermions in 2+1D at charge neutrality, the true ground state is not translationally invariant, but is a lattice of electron and hole filled regions. Tunneling transitions between these regions change the spectrum such that new Dirac nodes emerge parkNewGenerationMassless2008. In the weak coupling description the enhancement of the DC electrical conductivity is due to the new low energy states in these new Dirac nodes.
@@
For the thermal conductivity $\bar{\kappa}_{DC}$, shown as a function of temperature in Figure 1(c), we may conduct a similar analysis. At large temperatures — roughly $T\gtrsim 0.3G$ as can be deduced from the inset in Figure 1(c) — a hydrodynamical analysis ought to be valid, such that the thermal transport is fully coherent with a Drude-like response. This is what the numerical data indeed appears to reflect, with a decrease of thermal conductivity with increasing $T/G$ explainable by a thermal broadening of a Drude peak. However, a closer inspection of the AC thermal conductivity will reveal that this is not true. Before we explain there the underlying physics, we do observe that at large $T\gg G$ the result is perturbative in the lattice strength $\delta\mu$: The thermal conductivity scales as $\bar{\kappa}_{DC}\sim(\delta\mu)^{-2}$ suggesting again (incorrectly) a single decay rate $\Gamma\sim(\delta\mu)^{2}$ (cf. upper inset in Figure 1(d)). Increasing $\delta\mu$ therefore means decreasing $\bar{\kappa}$. However, it can be shown that the thermal conductivity is also bounded from below as $\kappa/T\geq 4\pi^{2}/3$ grozdanov_incoherent_2016, meaning that the thermal conductivity must eventually saturate. We indeed find that this is the case, as is shown in the lower inset in Figure 1(d), where for the largest lattice strengths that we are able to probe ($\delta\mu=20G$ at $T=0.1G$) the thermal conductivity stays above the bound by about a factor of $1.4$.
@@ 4.2 AC transport in 1D charge-neutral lattices
Next, we study the finite frequency linear response. In the high temperature regime, $T\gg G$}, $T\geq\delta\mu$} shown in Figure 3(b) and (c), the AC electrical conductivity $\sigma(\omega)$ is completely dominated by an Umklapped sound mode, which sits at $\omega=v_{\text{sound}}G\approx G/\sqrt{2}$ and which becomes increasingly sharp at larger $T/G$. The identification is readily made by comparing with systems at finite $\bar{\mu}$ donos_holographic_2014; balm_t-linear_2023. This may seem surprising, as in a charge-neutral system one expects electrical and thermal transport to decouple, with the sound mode residing entirely in the thermal sector. This decoupling between electrical and thermal transport is however only valid at leading order in the charge density. At subleading order (e.g., in terms of the density-density susceptibility $\chi_{nn}$), the cross-coupling between sound and charge diffusion does survive chagnet_hydrodynamics_2024.
@@
These two poles must have arisen from a collision of two diffusive poles constrained to the imaginary frequency axis. The realization that Umklapp sound appears in the electrical conductivity $\sigma(\omega)$ points to an answer for the origin of these two poles. In the presence of spatial periodic modulations, hydrodynamic excitations are not given by Fourier modes, but by Bloch waves. For weak interactions these can be interpreted as Umklapped versions of original hydrodynamic excitations. In the case of charged hydrodynamics, this Umklapp effect predicts that the single Drude mode gets accompanied by two Umklapp sound modes and one Umklapp charge diffusion mode chagnet_hydrodynamics_2024. Generically both electrical and AC thermal transport will be dominated by these four modes at low frequencies. However, as first observed in balm_t-linear_2023, for certain parameters the (diffusive) Drude mode and the Umklapp charge diffusion mode can collide and move symmetrically off the imaginary axis, acquiring a propagating part. For finite $\bar{\mu}$ this happens as $T/\bar{\mu}$ is increased. By adding a small overall chemical potential such that we’re back in the regime where both modes are diffusive, we see in Figure 5 that the same physics happens as we tune $\bar{\mu}/T$ to zero instead. This explains these two poles. Umklapp hydrodynamics predicts four poles, but as $\bar{\mu}\rightarrow 0$ the residue of the two other poles — Umklapp sound — becomes negligible in the AC thermal conductivity $\bar{\kappa}(\omega)$ (cf. Figure 4). This can be confirmed analytically chagnet_hydrodynamics_2024. Likewise, we see that at charge neutrality the residue of the Drude and charge diffusion pole in the AC electrical conductivity $\sigma(\omega)$ also vanishes. The set of four poles thus divide themselves between the electrical and thermal response at $\bar{\mu}=0$; Umklapp sound is purely in the electrical sector, whereas the diffusive poles (after the collision) govern the thermal response.
These exotic pole collisions near $\mu/T=1$ are by now well known to occur in strongly coupled systems; for instance, it is also responsible for a vanishing plasma frequency for $\bar{\mu}/T\lesssim 1$ in models with dynamical electromagnetism gran_exotic_2019. For any finite $\delta\mu$, the thermal response is thus never truly Drude: unlike the hydrodynamic expressions 25, the momentum relaxation rate does not enter through a single Drude pole, but always through this pair.
@@
At even lower temperatures $T=0.1G$ (Figure 3, panels (a) and (d)), the resonance peak in the electrical conductivity $\sigma(\omega)$ starts to move towards lower frequencies, becoming slightly more sharply peaked at higher $\delta\mu/G$ as compared to moderate temperatures. [^3] Note that the increase in DC electrical conductivity with stronger lattices (Figure 1) is barely visible in the $\omega\to 0$ limit in Figure 3 due to the scale. As is evident from Figure 4, this response is still dominated by the single pole that originates in Umklapp sound, though this regime is quite far from the Umklapp hydrodynamics regime $\delta\mu\ll T,\bar{\mu}$, and the imprint of the pole on the real axis is no longer close to $\omega=v_{\text{sound}}G$. This deformed Umklapp-sound peak remains absent in thermal transport: the physical mechanisms between thermal and electrical transport stay distinct at low temperatures. In the AC thermal conductivity $\bar{\kappa}(\omega)$, on the other hand, the mid-IR peak originating in the collision of the Drude pole with the Umklapp charge diffusion pole sharpens and moves towards $\omega=G$ as $T$ decreases. Though a peak at $\omega=G$ asks for an intuitive explanation in terms of the lattice and Umklapp, no such obvious explanation exists. Most such attempts will inadvertently rely on hydrodynamics, but for these lattice amplitudes one is outside its range of validity. A possible explanation is that this is a large $N$ artifact denoting a “graviton-exchange” in the holographic gravitational description as postulated in chesler_conformal_2014. Such a collective mode would travel at the speed of light, and can therefore give rise to an $\omega=G$ excitation after Umklapp. As this happens at $T<G$, an $\omega=G$ energy-mode is outside of the hydrodynamical regime, precluding an easy understanding from the quantum critical theory perspective.
@@
5 Thermoelectric transport in 2D charge neutral lattices

### v2

@@ 3 Transport expectations from hydrodynamics
The periodic potential environment in which we wish to study the quantum critical response, is captured in the UV boundary conditions of our ansatz 10. At finite density, i.e., finite average $\bar{\mu}\neq 0$, a simple (relativistic) hydrodynamic analysis extended with a momentum relaxation rate $\Gamma$ parametrically smaller than any other rate ($\bar{\mu}\gg\delta\mu$ in the case of an chemical potential lattice) results in thermo-electric transport that is dominated by a Drude pole at $\omega=-i\Gamma\sim-i(\delta\mu)^{2}$, such that the AC transport coefficients take the general form Davison and Goutéraux (2015); Hartnoll et al. (2007)
@@
where $n,s,\varepsilon$ and $p$ denote the charge, entropy, energy and pressure densities, respectively. We will be studying the system in the opposite limit $\bar{\mu}\ll\delta\mu$, but the generic wisdom that translational symmetry breaking leads to finite DC conductivities should hold. In the extremal limit of interest to us, $\bar{\mu}=n=0$, a simple exercise in hydrodynamics indicates that electric and thermal transport will decouple. We might therefore expect the electrical conductivity to be governed by the incoherent conductivity $\sigma_{Q}$. Similarly, the thermal conductivity should remain coherent (i.e., dominated by a single pole). In the absence of magnetic fields, we furthermore expect $\alpha=0$, confirming that the mechanisms of electrical and thermal transport are essentially independent. We thus also expect no difference between the thermal conductivity at zero electric current $\kappa_{ij}=\bar{\kappa}_{ij}-T\bar{\alpha}_{ik}(\sigma^{-1})_{kl}\alpha_{lj}$} and $\bar{\kappa}_{ij}$.
At finite magnetic fields, the transport coefficients become tensorial in nature, and one naturally expects $\sigma_{xy}\sim n/B$, as follows from a combination of parity and charge-conjugation symmetry. In a charge neutral system with $n=0$, however, no Hall effect will occur, and the magnetic field will only affect longitudinal transport. Nevertheless, a finite magnetic field does allow for a Hall thermoelectric conductivity $\alpha_{xy}\sim s/B$ and we will be able to observe a non-zero contribution to the Nernst coefficient $e_{N}=-(\sigma^{-1}\alpha)_{xy}$, where from charged hydrodynamics one expects Hartnoll et al. (2007)
@@
with the cyclotron frequency in natural units equal to $\omega_{c}=\frac{Bn}{\epsilon+p}$. [^4] A review of charged hydrodynamics in the presence of weak translational symmetry breaking is given in Amoretti and Brattan (2022). For a charge neutral system this simplifies to a direct measurement of the momentum relaxation rate as $e_{N}=\frac{B}{T\Gamma}$.
Figure 1: DC quantum critical transport in a 1D charge-neutral lattice, $\mu(x)=\delta\mu\cos(Gx)$}. Note that the thermoelectric conductivity $\alpha=0$} (numerically confirmed, not shown) and thus $\bar{\kappa}=\kappa$. (a) Electrical conductivity $\sigma$ as a function of $T/\delta\mu$ for various $\delta\mu/G$ (left colorbar). At low temperatures, the lattice causes the conductivity to increase. Inset: electrical conductivity as a function of $T/G$. (b) Electrical conductivity $\sigma$ as a function of lattice amplitude for various temperatures $T/G$ (right colorbar). In the low-temperature limit for small $\delta\mu/G$, the scaling is that predicted by Chesler, Lucas and Sachdev Chesler et al. (2014), (Eq. 28 black dashed line). Inset: difference from the CFT value $\sigma_{\infty}=1$ on a logarithmic scale, which grows as $(\delta\mu)^{2}$. (c) Thermal conductivity $\bar{\kappa}$ as a function of $T/\delta\mu$. There is a cross-over from a universal high-$T$ regime to a strong thermal conductor-regime at low $T$. Inset: thermal conductivity $\bar{\kappa}$ as a function of $T/G$. (d) Thermal conductivity $\bar{\kappa}$ as a function of lattice amplitude. The thermal conductivity decreases most rapidly with increased lattice amplitude around $T/G\approx 0.1$}, which corresponds with the minima seen in panel (c). Upper inset: for weak lattices, the thermal conductivity scales as $1/(\delta\mu)^{2}$. Lower inset: for strong lattices, the thermal conductivity approaches the bound $\kappa/T\geq 4\pi^{2}/3$ of Grozdanov et al. (2016).
There is one more expectation we can formulate. Our system is on average isotropic in the thermodynamic limit. In quantum critical theories (with a holographic dual), this implies that the DC electrical conductivity is bounded from below as $\sigma_{DC}\geq\sigma_{\infty}$ Grozdanov et al. (2015). The electrical conductivity at charge neutrality in the absence of a lattice is given by that of the CFT dual to the Schwarzchild AdS4 solution: $\sigma_{DC}=\sigma_{\infty}=1$. This saturates the bound. Thus, the only thing the conductivity can do is to increase with increased disorder/increased lattice amplitude around $\bar{\mu}=0$. We will indeed see this below.
@@ 4.1 DC transport in 1D charge-neutral lattices
To build on the results of Chesler, Lucas and Sachdev Chesler et al. (2014), we first consider a 1D lattice at charge neutrality (i.e., that is charge-conjugation symmetric): we set the background chemical potential to be
@@
This scaling of the electrical conductivity $\sigma_{DC}$ as a function of lattice strength was elucidated by Chesler, Lucas and Sachdev Chesler et al. (2014): deforming the UV with a 1D lattice with $\bar{\mu}=0$ still results in an AdS4 IR geometry at $T=0$, except that in the natural UV coordinates the IR metric is anisotropic with regards to the $x/y$-directions. We highlight this effect in Figure 2 (left column), that shows the RG flow (captured by the radial coordinate $z$), but halted due to the finite value of $T$, of the density vanishing in the IR and two metric components exhibiting the anisotropy. Absorbing the emergent anisotropy in the coordinates to bring the IR metric into standard AdS4 form also rescales the length scales of the dual theory, resulting in
@@
This perturbative result, valid for $T=0$ and at $\delta\mu\ll G$, is shown as a black dashed line in Figure 1(b) and which can be seen to asymptotically hold for $T\ll G$; in fact, the result extends to $\delta\mu/G=O(1)$ for low enough temperatures. At larger temperatures the deviation from $\sigma_{\infty}$ is not as strong, but still scales as $(\delta\mu)^{2}$ (inset, panel (b)). Two comments about this strong coordinate anisotropy hiding translational symmetry mechanism are in order. Firstly, the explanation is particular to 1D, and will not apply to the 2D lattice as we shall see in Section 5. Secondly, as explained by Chesler, Lucas and Sachdev, this emergent translational symmetry is an artifact of the holographic large $N$ approximation Chesler et al. (2014). There it was shown that by comparing to a system of $N_{f}$ free Dirac fermions in 2+1D at charge neutrality, the true ground state is not translationally invariant, but is a lattice of electron and hole filled regions. Tunneling transitions between these regions change the spectrum such that new Dirac nodes emerge Park et al. (2008b). In the weak coupling description the enhancement of the DC electrical conductivity is due to the new low energy states in these new Dirac nodes.
@@
For the thermal conductivity $\bar{\kappa}_{DC}$, shown as a function of temperature in Figure 1(c), we may conduct a similar analysis. At large temperatures — roughly $T\gtrsim 0.3G$ as can be deduced from the inset in Figure 1(c) — a hydrodynamical analysis ought to be valid, such that the thermal transport is fully coherent with a Drude-like response. This is what the numerical data indeed appears to reflect, with a decrease of thermal conductivity with increasing $T/G$ explainable by a thermal broadening of a Drude peak. However, a closer inspection of the AC thermal conductivity will reveal that this is not true. Before we explain there the underlying physics, we do observe that at large $T\gg G$ the result is perturbative in the lattice strength $\delta\mu$: The thermal conductivity scales as $\bar{\kappa}_{DC}\sim(\delta\mu)^{-2}$ suggesting again (incorrectly) a single decay rate $\Gamma\sim(\delta\mu)^{2}$ (cf. upper inset in Figure 1(d)). Increasing $\delta\mu$ therefore means decreasing $\bar{\kappa}$. However, it can be shown that the thermal conductivity is also bounded from below as $\kappa/T\geq 4\pi^{2}/3$ Grozdanov et al. (2016), meaning that the thermal conductivity must eventually saturate. We indeed find that this is the case, as is shown in the lower inset in Figure 1(d), where for the largest lattice strengths that we are able to probe ($\delta\mu=20G$ at $T=0.1G$) the thermal conductivity stays above the bound by about a factor of $1.4$.
@@ 4.2 AC transport in 1D charge-neutral lattices
Next, we study the finite frequency linear response. In the high temperature regime, $T\gg G$}, $T\geq\delta\mu$} shown in Figure 3(b) and (c), the AC electrical conductivity $\sigma(\omega)$ is completely dominated by an Umklapped sound mode, which sits at $\omega=v_{\text{sound}}G\approx G/\sqrt{2}$ and which becomes increasingly sharp at larger $T/G$. The identification is readily made by comparing with systems at finite $\bar{\mu}$ Donos and Gauntlett (2014a); Balm et al. (2023). This may seem surprising, as in a charge-neutral system one expects electrical and thermal transport to decouple, with the sound mode residing entirely in the thermal sector. This decoupling between electrical and thermal transport is however only valid at leading order in the charge density. At subleading order (e.g., in terms of the density-density susceptibility $\chi_{nn}$), the cross-coupling between sound and charge diffusion does survive Chagnet and Schalm (2024).
@@
These two poles must have arisen from a collision of two diffusive poles constrained to the imaginary frequency axis. The realization that Umklapp sound appears in the electrical conductivity $\sigma(\omega)$ points to an answer for the origin of these two poles. In the presence of spatial periodic modulations, hydrodynamic excitations are not given by Fourier modes, but by Bloch waves. For weak interactions these can be interpreted as Umklapped versions of original hydrodynamic excitations. In the case of charged hydrodynamics, this Umklapp effect predicts that the single Drude mode gets accompanied by two Umklapp sound modes and one Umklapp charge diffusion mode Chagnet and Schalm (2024). Generically, both electrical and AC thermal transport will be dominated by these four modes at low frequencies. However, as first observed in Balm et al. (2023), for certain parameters the (diffusive) Drude mode and the Umklapp charge diffusion mode can collide and move symmetrically off the imaginary axis, acquiring a propagating part. For finite $\bar{\mu}$ this happens as $T/\bar{\mu}$ is increased. By adding a small overall chemical potential such that we’re back in the regime where both modes are diffusive, we see in Figure 5 that the same physics happens as we tune $\bar{\mu}/T$ to zero instead. This explains these two poles. Umklapp hydrodynamics predicts four poles, but as $\bar{\mu}\rightarrow 0$ the residue of the two other poles — Umklapp sound — becomes negligible in the AC thermal conductivity $\bar{\kappa}(\omega)$ (cf. Figure 4). This can be confirmed analytically Chagnet and Schalm (2024). Likewise, we see that at charge neutrality the residue of the Drude and charge diffusion pole in the AC electrical conductivity $\sigma(\omega)$ also vanishes. The set of four poles thus divide themselves between the electrical and thermal response at $\bar{\mu}=0$; Umklapp sound is purely in the electrical sector, whereas the diffusive poles (after the collision) govern the thermal response.
These exotic pole collisions near $\bar{\mu}/T=1$ are by now well known to occur in strongly coupled systems; for instance, it is also responsible for a vanishing plasma frequency for $\bar{\mu}/T\lesssim 1$ in models with dynamical electromagnetism Gran et al. (2019). For any finite $\delta\mu$, the thermal response is thus never truly Drude: unlike the hydrodynamic expressions 25, the momentum relaxation rate does not enter through a single Drude pole, but always through this pair.
@@
At even lower temperatures $T=0.1G$ (Figure 3, panels (a) and (d)), the resonance peak in the electrical conductivity $\sigma(\omega)$ starts to move towards lower frequencies, becoming slightly more sharply peaked at higher $\delta\mu/G$ as compared to moderate temperatures. [^5] Note that the increase in DC electrical conductivity with stronger lattices (Figure 1) is barely visible in the $\omega\to 0$ limit in Figure 3 due to the scale. As is evident from Figure 4, this response is still dominated by the single pole that originates in Umklapp sound, though this regime is quite far from the Umklapp hydrodynamics regime $\delta\mu\ll T,\bar{\mu}$, and the imprint of the pole on the real axis is no longer close to $\omega=v_{\text{sound}}G$. This deformed Umklapp-sound peak remains absent in thermal transport: the physical mechanisms between thermal and electrical transport stay distinct at low temperatures. In the AC thermal conductivity $\bar{\kappa}(\omega)$, on the other hand, the mid-IR peak originating in the collision of the Drude pole with the Umklapp charge diffusion pole sharpens and moves towards $\omega=G$ as $T$ decreases. Though a peak at $\omega=G$ asks for an intuitive explanation in terms of the lattice and Umklapp, no such obvious explanation exists. Most such attempts will inadvertently rely on hydrodynamics, but for these lattice amplitudes one is outside its range of validity. A possible explanation is that this is a large $N$ artifact denoting a “graviton-exchange” in the holographic gravitational description as postulated in Chesler et al. (2014). Such a collective mode would travel at the speed of light, and can therefore give rise to an $\omega=G$ excitation after Umklapp. As this happens at $T<G$, an $\omega=G$ energy-mode is outside of the hydrodynamical regime, precluding an easy understanding from the quantum critical theory perspective.
@@
5 Thermoelectric transport in 2D charge-neutral lattices